OpenStudy (anonymous):

Find the area of each rectangle with the base and height provided 30 in, 4 yd

OpenStudy (anonymous):

I'd figure the area in similar units--I'm going to use inches 4 yards = 36*4 = 144 inches Area of a triangle is 1/2bh=1/2(30)(144)=2160 sq inches